# Invariants of Finite Orthogonal Groups of Plus Type in Odd Characteristic

### v2

@@ 5.1. Computing the Hook Invariants
We work in the ring $\widetilde{T}:=T[a_{1},a_{2},\dots,a_{q}]$ where $a_{1},a_{2},\dots,a_{q}$ are $q$ new indeterminants. Write $f=\sum_{k}\sum_{j}c_{k,j}y_{1}^{k}t^{j}$. We have $\nu(c_{k,j})\geq||k||-1$ for all $k$ and $j$. Consider
$\widetilde{f}$ $=\prod_{s=1}^{q}f(y_{1}+a_{s}u,t-a_{s}x_{1})\in\widetilde{T}$
$=\prod_{s=1}^{q}\left(\sum_{k}\sum_{j}c_{k,j}(y_{1}+a_{s}u)^{k}(t-a_{s}x_{1})^{j}\right)$
@@
Clearly $\widetilde{f}$ is invariant under any permutation of the $a_{s}$. Thus we may express $\widetilde{f}$ using the elementary symmetric functions $E_{i}(a_{1},a_{2},\dots,a_{q})$ of degree $i$ in the $a_{i}$, i.e., $\widetilde{f}=\sum_{\vec{e}}h_{(e_{1},e_{2},\dots,e_{q})}E_{1}^{e_{1}}E_{2}^{e_{2}}\cdots E_{q}^{e_{q}}$ with $h_{(e_{1},e_{2},\dots,e_{q})}\in T$. When we specialize the $a_{i}$ to the elements of the field $\mathbb{F}_{q}=\{\alpha_{1},\alpha_{2},\dots,\alpha_{q}\}$, $\widetilde{f}$ specializes to $\mathcal{T}_{(y_{1},u,t,x_{1})}(f)$.
Note that $E_{i}(\alpha_{1},\alpha_{2},\dots,\alpha_{q})=0$ for $i\neq q-1$ and $i\geq 1$. This follows from
$X^{q}-X=\prod_{a\in\mathbb{F}_{q}}(X-a)=\sum_{i=0}^{q}(-1)^{i}E_{i}(\alpha_{1},\alpha_{2},\dots,\alpha_{q})X^{q-i}$
@@
Terms occurring in $h_{(0,0,\dots,0,e,0)}$ satisfy $i+\ell=e(q-1)$. The term $h_{(0,0,\dots,0,0,0)}$ when $i=\ell=0$ is just $h_{(0,0,\dots,0,0,0)}=\prod_{s=1}^{q}f(y_{1},t)=f(y_{1},t)^{q}$. Since $||k^{q}||=||k||$ it is clear that $h_{(0,0,\dots,0,0,0)}$ is compliant.
@@
For $h_{(0,0,\dots,0,e,0)}$ with $e\geq 1$ we have $i+\ell=e(q-1)$ is a positive multiple of $q-1$. Thus $||i+\ell||$ is also a positive multiple of $q-1$ by Lemma 5.6. In particular $||i+\ell||\geq q-1$.
Thus
$||k||=||\sum_{s=1}^{q}k_{s}-i_{s}||\leq\sum_{s=1}^{q}||k_{s}-i_{s}||=\sum_{s=1}^{q}(||k_{s}||-||i_{s}||)\ .$
@@
$\ell+\sum_{s=1}^{q}||k_{s}||$ $\geq\ell+||k||+\sum_{s=1}^{q}||i_{s}||\geq\ell+||k||+||\sum_{s=1}^{q}i_{s}||=\ell+||k||+||i||$
$\geq||\ell||+||k||+||i||\geq||k||+||i+\ell||\geq||k||+(q-1)$
Thus $\nu(u^{i}t^{j}x_{1}^{\ell}\prod_{s=1}^{q}c_{k_{s},j_{s}})\geq\ell+\sum_{s=1}^{q}(||k_{s}||-1)\geq||k||-1$ as required. ∎
Let $\overline{H}$ denote the pointwise stabiliser of $\{f_{m},e_{m}\}$ in $H$ and observe that $\overline{H}$ also stabilises $\{y_{m},x_{m}\}$. It is easy to see that $\overline{H}$ is a normal subgroup of $H$. Furthermore, we can identify $H/\overline{H}$ with the subgroup of $H$ consisting of elements which stabilise all variables except $y_{1}$, $y_{m}$, and $x_{m}$. Observe that
$N(y_{1})=\prod_{a,b\in\mathbb{F}_{q}}N^{\overline{H}}(y_{1}+by_{m}+ax_{m}-abx_{1})$
where $N^{\overline{H}}(v)$ denotes the orbit product of $v$ over $\overline{H}$. By the induction hypothesis, $N^{\overline{H}}(y_{1})$ is compliant. Since $\deg_{x_{m}}(N^{\overline{H}}(y_{1}))=0$, we can apply Lemma 5.7 to $N^{\overline{H}}(y_{1})$ with $u=t=x_{m}$ to show that
$M(y_{1},x_{m}):=\mathcal{T}_{(y_{1},x_{m},x_{m},x_{1})}(N^{\overline{H}}(y_{1}))=\prod_{a\in\mathbb{F}_{q}}N^{\overline{H}}(y_{1}+ax_{m})$
is compliant. Applying Lemma 5.7 to $M(y_{1},x_{m})$ with $u=y_{m}$, we see that
$\mathcal{T}_{(y_{1},y_{m},x_{m},x_{1})}(M)$ $=\prod_{b\in\mathbb{F}_{q}}M(y_{1}+by_{m},x_{m}-bx_{1})$
$=\prod_{b\in\mathbb{F}_{q}}\bigg{(}\prod_{a\in\mathbb{F}_{q}}N^{\overline{H}}(y_{1}+by_{m}+a(x_{m}-bx_{1})\bigg{)}$
$=\prod_{b\in\mathbb{F}_{q}}\prod_{a\in\mathbb{F}_{q}}N^{\overline{H}}(y_{1}+by_{m}+ax_{m}-abx_{1})$
$=N(y_{1})$
is compliant.
Since $N(y_{1})$ is compliant, $x_{1}N(y_{1})$ is strongly compliant. Thus $x_{1}N(y_{1})-\xi_{n-2}$ is strongly compliant. Since $\deg_{y_{1}}(x_{1}N(y_{1})-\xi_{n-2})<q^{n-2}$, using Lemma 5.5, we see that
$x_{1}N(y_{1})-\xi_{n-2}\in Q^{H}[\xi_{i}\mid n=0,\ldots,n-3].$
@@ 5.8. Computing the Sylow Invariants
Recall that $H$ is a normal subgroup of $P_{m}$. Further more, identifying $P_{m-1}$ with pointwise stabiliser of $\{x_{1},y_{1}\}$ in $P_{m}$, we have $P_{m}=P_{m-1}H$. Therefore $S_{m}^{P_{m}}=(S_{m}^{H})^{P_{m-1}}$. Let $\mathcal{B}_{H}$ denote the monomial factors of $\prod_{i=0}^{n-3}\xi_{i}^{q-1}$. Using the results of Section 5.1, we see that $S_{m}^{H}$ is the free module over $Q^{H}[N(y_{1})]$ with basis given by the elements of $\mathcal{B}_{H}$. Thus
$S_{m}^{H}=\bigoplus_{\gamma\in\mathcal{B}_{H}}Q^{H}[N(y_{1})]\gamma.$
@@
$S_{m}^{P_{m}}=(S_{m}^{H})^{P_{m-1}}=\bigoplus_{\gamma\in\mathcal{B}_{H}}(Q^{H}[N(y_{1})])^{P_{m-1}}\gamma.$
Write $\overline{S}$ for the subalgebra of $Q^{H}$ constructed by omitting the generator $x_{1}$. Then $Q^{H}[N(y_{1})]=\overline{S}\otimes\mathbb{F}_{q}[x_{1},N(y_{1})]$ and
$(Q^{H}[N(y_{1})])^{P_{m-1}}=\overline{S}^{P_{m-1}}\otimes\mathbb{F}_{q}[x_{1},N(y_{1})].$
The action of $P_{m-1}$ on $\overline{S}$ is equivalent to the action of $P_{m-1}$ on $S_{m-1}$. This gives an algebra isomorphism from $S_{m-1}^{P_{m-1}}$ to $\overline{S}^{P_{m-1}}$ which takes elements of degree $d$ to elements of degree $qd$. This isomorphism is induced by $\varphi:=\psi_{1}\circ\sigma$, where $\sigma$ is the algebra monomorphism from $S_{m-1}$ to $S_{m}$ which takes $y_{i}$ to $y_{i+1}$, $x_{i}$ to $x_{i+1}$. Note that the isomorphism takes $\xi_{i}\in S_{m-1}$ to $\psi_{1}(\xi_{i})\in S_{m}$.
@@
For the induction step, first observe that $S_{m}^{P_{m}}$ is generated by $x_{1}$, $N(y_{1})$, elements of $\mathcal{B}_{H}$ and generators for $\overline{S}^{P_{m-1}}$. The isomorphism from $S_{m-1}^{P_{m-1}}$ to $\overline{S}^{P_{m-1}}$ takes orbit products to orbit products and $\xi_{i}$ to $\psi_{1}(\xi_{i})\in\mathbb{F}_{q}[x_{1},\xi_{j}\mid j\leq i+1]$. Thus $S_{m}^{P_{m}}$ is generated by orbit products and the $\xi_{i}$. By induction, $S_{m-1}^{P_{m-1}}$ is a free module over the algebra generated by the orbit products of the variables. Let $\mathcal{C}$ denote the image of the module generators in $\overline{S}^{P_{m-1}}$. Then $\mathcal{B}_{H}\mathcal{C}$ is the set of module generators for $S_{m}^{P_{m}}$ over the algebra generated by the orbit products of the variables and $S_{m}^{P_{m}}$ is Cohen-Macaulay. ∎
Define algebra maps $\psi_{j}:S_{m}\to S_{m}$ by
$\psi_{j}(v):=\prod_{u\in U_{j}}(v-u)$
for $v$ homogeneous of degree one, where $U_{j}$ is the span of $\{x_{1},\ldots,x_{j}\}$. These maps $\psi_{j}$ were introduced and used effectively by Ferreira and Fleischmann in [14].
@@
In the lexicographic term order $\operatorname{LT}(\psi_{j}(\xi_{i}))=y_{j+1}^{q^{i+j}}x_{j+1}^{q^{j}}$.
@@
The linear forms $x_{1},x_{2},\dots,x_{j}$ lie in the kernel of $\psi_{j}$. Furthermore for $k>j$ and $a,b$ positive integers $\operatorname{LT}(\psi_{j}(y_{k}^{a}x_{k}^{b}))=y_{k}^{q^{j}a}x_{k}^{q^{j}b}$. From these observations the result follows. ∎
@@
$\varphi\circ\psi_{j}=\psi_{j+1}\circ\sigma$ and $\varphi(\psi_{j}(\xi_{i}))=\psi_{j+1}(\xi_{i})$.
@@
For $j=0$, the first equation is the definition of $\varphi$ and the second follows from the observation that $x_{1}$ is in the kernel of $\psi_{1}$.
Suppose $j>0$. Both $\varphi\circ\psi_{j}$ and $\psi_{j+1}\circ\sigma$ are algebra maps so to prove the first equation it is sufficient to show that $\varphi(\psi_{j}(v))=\psi_{j+1}(\sigma(v))$ when $\deg(v)=1$. We have
$\varphi(\psi_{j}(v))=\psi_{1}\left(\sigma\left(\prod_{u\in U_{j}}(v-u)\right)\right)=\psi_{1}\left(\prod_{u\in U_{j}}(\sigma(v)-\sigma(u))\right).$
However, $U_{j+1}=U_{1}\oplus\sigma(U_{j})$. Therefore
$\psi_{1}\left(\prod_{u\in U_{j}}(\sigma(v)-\sigma(u))\right)=\prod_{u\in U_{j+1}}(\sigma(v)-u)=\psi_{j+1}(\sigma(v))$
as required. For the second equation, we have $\varphi(\psi_{j}(\xi_{i}))=\psi_{j+1}(\sigma(\xi_{i}))=\psi_{j+1}(\xi_{i})$, since $x_{1}$ is in the kernel of $\psi_{j+1}$. ∎
@@
In this example we consider the Sylow invariants for ${{\text{O}^{+}_{{6}}(\mathbb{F}_{q})}}$. It follows from Theorem 5.9 that $S_{3}^{P_{3}}$ is generated by the homogeneous system of parameters $\mathcal{H}_{0}:=\{N(y_{1}),N(y_{2}),N(y_{3}),N(x_{3}),N(x_{2}),x_{1}\}$ and $\{\xi_{0},\xi_{1},\xi_{2},\xi_{3}\}$. Using the lexicographic order the lead monomials are
$y_{1}^{q^{4}},y_{2}^{q^{3}},y_{3}^{q^{2}},x_{3}^{q^{2}},x_{2}^{q},x_{1},y_{1}x_{1},y_{1}^{q}x_{1},y_{1}^{q^{2}}x_{1},y_{1}^{q^{3}}x_{1}.$

### v3

@@ 5.1. Computing the Hook Invariants
[…]
$\widetilde{f}$ $=\prod_{s=1}^{q}f(y_{m}+\alpha_{s}t,x_{1}-\alpha_{s}x_{m})\in\widetilde{T}$
$=\prod_{s=1}^{q}\left(\sum_{k}\sum_{j}r_{k,j}(y_{m}+\alpha_{s}t)^{k}(x_{1}-\alpha_{s}x_{m})^{j}\right)$
@@
Clearly $\widetilde{f}$ is invariant under any permutation of the $\alpha_{s}$. Thus we may express $\widetilde{f}$ using the elementary symmetric functions $E_{i}(\alpha_{1},\alpha_{2},\dots,\alpha_{q})$ of degree $i$ in the $\alpha_{i}$, i.e., $\widetilde{f}=\sum_{\vec{e}}h_{(e_{1},e_{2},\dots,e_{q})}E_{1}^{e_{1}}E_{2}^{e_{2}}\cdots E_{q}^{e_{q}}$ with $h_{(e_{1},e_{2},\dots,e_{q})}\in T$. When we specialize the elementary symmetric functions $E_{i}(\alpha_{1},\alpha_{2},\dots,\alpha_{q})$ to the values $E_{i}(\mathbb{F}_{q})$, $\widetilde{f}$ specializes to $\mathcal{T}_{t}(f)$.
Note that $E_{i}(\mathbb{F}_{q})=0$ for $i\neq q-1$ and $i\geq 1$. This follows from
$X^{q}-X=\prod_{a\in\mathbb{F}_{q}}(X-a)=\sum_{i=0}^{q}(-1)^{i}E_{i}(\mathbb{F}_{q})X^{q-i}.$
@@
Terms occurring in $h_{(0,0,\dots,0,e,0)}$ satisfy $i+\ell=e(q-1)$. The term $h_{(0,0,\dots,0,0,0)}$ when $i=\ell=0$ is just $h_{(0,0,\dots,0,0,0)}=\prod_{s=1}^{q}f(y_{m},x_{1})=f(y_{m},x_{1})^{q}$. Since kq=k||kq||=||k|| it is clear that $h_{(0,0,\dots,0,0,0)}$ is $(y_{m},x_{m})$-compliant.
@@
[…]
For $h_{(0,0,\dots,0,e,0)}$ with $e\geq 1$ we have $i+\ell=e(q-1)$ is a positive multiple of q1q-1. Thus $i+\ell||$ is also a positive multiple of q1q-1 by Lemma 5.8. In particular $i+\ell||\geq q-1$.
@@
[…]
Let $\overline{H}$ denote the pointwise stabiliser of $\{f_{1},e_{1}\}$ in $H=H_{m}$ and observe that $\overline{H}$ also stabilises $\{y_{1},x_{1}\}$. Using a relabeling of the variables which takes $y_{i}$ to $y_{i+1}$and $x_{i}$ to $x_{i+1}$, we can identify $H_{m-1}$ with $\overline{H}$. By induction $N^{H_{m-1}}(y_{m-1})$ is $(y_{m-1},x_{m-1})$-compliant. Using the identification of $H_{m-1}$ and $\overline{H}$, this implies that $N^{\overline{H}}(y_{m})$ is $(y_{m},x_{m})$-compliant (we use $N^{\overline{H}}(a)$ to denote the orbit product of $a$ over $\overline{H}$).
It is easy to see that $\overline{H}$ is a normal subgroup of $H$. Furthermore, we can identify $H/\overline{H}$ with the subgroup of $H$ consisting of elements which stabilise all variables except $y_{1}$, $y_{m}$, and $x_{1}$. Observe that
$N(y_{m})=\prod_{a,b\in\mathbb{F}_{q}}N^{\overline{H}}(y_{m}+by_{1}+ax_{1}-abx_{m})$
Since $N^{\overline{H}}(y_{m})$ is $(y_{m},x_{m})$-compliant and $\deg_{x_{1}}(N^{\overline{H}}(y_{m}))=0$, we can apply Lemma 5.9 to $N^{\overline{H}}(y_{m})$ with $t=x_{1}$ to show that
$M(y_{m},x_{1}):=\mathcal{T}_{x_{1}}(N^{\overline{H}}(y_{m}))=\prod_{a\in\mathbb{F}_{q}}N^{\overline{H}}(y_{m}+ax_{1})$
is $(y_{m},x_{m})$-compliant. Applying Lemma 5.9 to $M(y_{m},x_{1})$ with $t=y_{1}$, we see that
$\mathcal{T}_{y_{1}}(M)$ $=\prod_{b\in\mathbb{F}_{q}}M(y_{m}+by_{1},x_{1}-bx_{m})$
$=\prod_{b\in\mathbb{F}_{q}}\bigg(\prod_{a\in\mathbb{F}_{q}}N^{\overline{H}}(y_{m}+by_{1}+a(x_{1}-bx_{m})\bigg)$
$=\prod_{b\in\mathbb{F}_{q}}\prod_{a\in\mathbb{F}_{q}}N^{\overline{H}}(y_{m}+by_{1}+ax_{1}-abx_{m})$
$=N(y_{m})$
is $(y_{m},x_{m})$-compliant.
Since $N(y_{m})$ is $(y_{m},x_{m})$-compliant, $x_{m}N(y_{m})$ is strongly $(y_{m},x_{m})$-compliant. Thus $x_{m}N(y_{m})-\xi_{n-1}$ is strongly $(y_{m},x_{m})$-compliant. Since $\deg_{y_{m}}(x_{m}N(y_{m})-\xi_{n-1})<q^{n-2}$, using Lemma 5.7, we see that
$x_{m}N(y_{m})-\xi_{n-1}\in Q^{H}[\xi_{1},\dots,\xi_{n-2}].$
@@ 5.10. Computing the Sylow Invariants
Recall that $H$ is a normal subgroup of $P_{m}$. Furthermore, identifying $P_{m-1}$ with pointwise stabiliser of $\{x_{m},y_{m}\}$ in $P_{m}$, we have $P_{m}=P_{m-1}H$. Therefore $S_{m}^{P_{m}}=(S_{m}^{H})^{P_{m-1}}$. Let $\mathcal{B}_{H}$ denote the monomial factors of $\prod_{i=1}^{n-2}\xi_{i}^{q-1}$. Using the results of Section 5.1, we see that $S_{m}^{H}$ is the free module over $Q^{H}[N(y_{m})]$ with basis given by the elements of $\mathcal{B}_{H}$. Thus
$S_{m}^{H}=\bigoplus_{\gamma\in\mathcal{B}_{H}}Q^{H}[N(y_{m})]\gamma.$
@@
$S_{m}^{P_{m}}=(S_{m}^{H})^{P_{m-1}}=\bigoplus_{\gamma\in\mathcal{B}_{H}}(Q^{H}[N(y_{m})])^{P_{m-1}}\gamma.$
Write $\overline{S}$ for the subalgebra of $Q^{H}$ constructed by omitting the generator $x_{m}$. Then $Q^{H}[N(y_{m})]=\overline{S}\otimes\mathbb{F}_{q}[x_{m},N(y_{m})]$ and
$(Q^{H}[N(y_{m})])^{P_{m-1}}=\overline{S}^{P_{m-1}}\otimes\mathbb{F}_{q}[x_{m},N(y_{m})].$
The action of $P_{m-1}$ on $\overline{S}$ is equivalent to the action of $P_{m-1}$ on $S_{m-1}$. This gives an algebra isomorphism from $S_{m-1}^{P_{m-1}}$ to $\overline{S}^{P_{m-1}}$ which takes elements of degree $d$ to elements of degree $qd$. This isomorphism is the restriction of $\psi_{[m,1]}$ to $S_{m-1}^{P_{m-1}}\subset S_{m}$.
@@
For the induction step, first observe that $S_{m}^{P_{m}}$ is generated by $x_{m}$, $N(y_{m})$, elements of $\mathcal{B}_{H}$ and generators for $\overline{S}^{P_{m-1}}$. The isomorphism from $S_{m-1}^{P_{m-1}}$ to $\overline{S}^{P_{m-1}}$ takes orbit products to orbit products and $\xi_{i}$ to $\psi_{[m,1]}(\xi_{i})\in\mathbb{F}_{q}[x_{m},\xi_{j}\mid j\leq i+1]$. Thus $S_{m}^{P_{m}}$ is generated by orbit products and the $\xi_{i}$. By induction, $S_{m-1}^{P_{m-1}}$ is a free module over the algebra generated by the orbit products of the variables. Let $\mathcal{C}$ denote the image of the module generators in $\overline{S}^{P_{m-1}}$. Then $\mathcal{B}_{H}\mathcal{C}$ is the set of module generators for $S_{m}^{P_{m}}$ over the algebra generated by the orbit products of the variables and $S_{m}^{P_{m}}$ is Cohen-Macaulay. ∎
Recall from Section 2 that $\psi_{[m,j]}:S_{m}\to S_{m}$ is the algebra map defined by
$\psi_{[m,j]}(a):=\prod_{u\in U_{j}}(a-u)$
for $\deg(a)=1$, where $U_{j}=\operatorname{Span}_{\mathbb{F}_{q}}\{x_{m-j+1},\ldots,x_{m}\}$.
@@
In the lexicographic term order $\operatorname{LT}(\psi_{[m,j]}(\xi_{i+1}))=y_{m-j}^{q^{i+j}}x_{m-j}^{q^{j}}$.
@@
The linear forms $x_{m},x_{m-1},\dots,x_{m-j+1}$ lie in the kernel of $\psi_{[m,j]}$. Furthermore for $k\leq m-j$ and $a,b$ positive integers $\operatorname{LT}(\psi_{[m,j]}(y_{k}^{a}x_{k}^{b}))=y_{k}^{q^{j}a}x_{k}^{q^{j}b}$. From these observations the result follows. ∎
Let $\sigma$ denote the inclusion of $S_{m-1}$ into $S_{m}$. Define $\varphi:=\psi_{[m,1]}\circ\sigma$. The restriction of $\varphi$ gives the algebra isomorphism from $S_{m-1}^{P_{m-1}}$ to $\overline{S}^{{P_{m-1}}}\subset S_{m}$.
@@
$\varphi\circ\psi_{[m-1,j]}=\psi_{[m,j+1]}\circ\sigma$ and $\varphi(\psi_{[m-1,j]}(\xi_{i}))=\psi_{[m,j+1]}(\xi_{i})$.
@@
By definition, $\varphi=\psi_{[m,1]}\circ\sigma$. Since $x_{m}$ is in the kernel of $\psi_{[m,1]}$, we have $\varphi(\xi_{i})=\psi_{[m,1]}(\xi_{i})$. This proves the result for $j=0$, using the convention that $\psi_{[m-1,0]}$ is the identity map.
Suppose $j>0$. Both $\varphi\circ\psi_{[m-1,j]}$ and $\psi_{[m,j+1]}\circ\sigma$ are algebra maps from $S_{m-1}$ to $S_{m}$ so to prove the first equation it is sufficient to show that $\varphi(\psi_{[m-1,j]}(a))=\psi_{[m,j+1]}(\sigma(a))$ when $a$ is homogeneous of degree one. Define
$\widetilde{U}_{j}:=\operatorname{Span}_{\mathbb{F}_{q}}\{x_{m-j},\ldots,x_{m-1}\}\ .$
Then
$\varphi(\psi_{[m-1,j]}(a))=\psi_{[m,1]}\left(\sigma\left(\prod_{u\in\widetilde{U}_{j}}(a-u)\right)\right)=\psi_{[m,1]}\left(\prod_{u\in\widetilde{U}_{j}}(\sigma(a)-u)\right).$
Since $U_{j+1}=U_{1}\oplus\widetilde{U}_{j}$, we have
$\psi_{[m,1]}\left(\prod_{u\in\widetilde{U}_{j}}(\sigma(a)-u)\right)=\prod_{u\in U_{j+1}}(\sigma(a)-u)=\psi_{[m,j+1]}(\sigma(a))$
as required. For the second equation, since $x_{m}$ is in the kernel of $\psi_{[m,j+1]}$, we have $\psi_{[m,j+1]}(\xi_{i})=\psi_{[m,j+1]}(\sigma(\xi_{i}))=\varphi(\psi_{[m-1,j]}(\xi_{i}))$. ∎
@@
In this example we consider the Sylow invariants for $\text{O}^{+}_{{6}}(\mathbb{F}_{q})$}}. It follows from Theorem 5.11 that $S_{3}^{P_{3}}$ is generated by the homogeneous system of parameters $\mathcal{H}_{0}:=\{N(y_{3}),N(y_{2}),N(y_{1}),N(x_{1}),N(x_{2}),x_{3}\}$ and $\{\xi_{1},\xi_{2},\xi_{3},\xi_{4}\}$. Using the lexicographic order the lead monomials are
$y_{3}^{q^{4}},y_{2}^{q^{3}},y_{1}^{q^{2}},x_{1}^{q^{2}},x_{2}^{q},x_{3},y_{3}x_{3},y_{3}^{q}x_{3},y_{3}^{q^{2}}x_{3},y_{3}^{q^{3}}x_{3}.$